write the equation of the line in slope intercept form perpendicular to y=2/3x+6 through (6,5)
Answer by clarkp(14)

hello!!!
with the given equation which is y=2/3x+6 identify first the slope.the slope of that equation is....m= 2/3.
By then you had already identify use the formula "point slope form"
y-y(sub1)=m(x-xsub1).
now substitute. the pints that was give is (6,5)
y-5=2/3(x-6)
y-5=2/3x-12/3
y-5=(2/3)x-4 multiply both sides by 3 .
3(y-5)=3(2/3)x-4)3now distribute.
3y-15=2x-12 make it to general form.
-2x+3y-15+12=0
-2x+3y-3=0, in equation A must not be negative. To make the A positive ,multiply negative 1.
so the result the equation of the line is 2x-3y+3=0.